Hasan can do a job twice as fast as Jawed. Sameer can do a job as fast as Hasan and Jawed working together. If a certain job is completed in 4.5 hours by Jawed, how long it would take sameer to complete the same job?
A. 0.5 hour
B. 1 Hour
C. 1.5 Hours
D. 2 Hours

Answer by stanbon(75887)

Hasan can do a job twice as fast as Jawed. Sameer can do a job as fast as Hasan and Jawed working together. If a certain job is completed in 4.5 hours by Jawed, how long it would take sameer to complete the same job?
-----
Hasan DATA:
time = x hrs/job ; rate = 1/x job/hr
Jawed DATA:
time = 2x hrs/job ; rate = 1/(2x) job/hr
----
Sameer DATA:
rate = (1/x)+(1/2x) = (2x+x)/2x^2 = 3x/2x^2 = 3/(2x)
----------------------------------------------------
Equation:
If a certain job is completed in 4.5 hours by Jawed
4.5(1/(2x)) = 1 job
2x = 4.5
x = 2.25
------
Sameer time = 1/rate = (2x/3) = 2(2.25)/3 = 4.5/3 = 1.5 hrs
Cheers,
Stan H.
-------------
A. 0.5 hour
B. 1 Hour
C. 1.5 Hours
D. 2 Hours